OpenStudy (anonymous):

4x + 5y + 7(x – y) – 8x A little help please? :)

OpenStudy (anonymous):

are these inequalities or what category of math is this?

OpenStudy (anonymous):

There combine like terms in the distributive property

OpenStudy (triciaal):

are you suppose to simplify the above? what is the question?

OpenStudy (triciaal):

first distribute multiply each term inside the parenthesis by the term outside then group like terms to simplify put all the x's together, all the y's together all the numbers etc.

OpenStudy (anonymous):

simplify the combine like terms. This isn't my actual home work problem, I switched the numbers

OpenStudy (triciaal):

7(x-y) = 7*x -7*y

OpenStudy (triciaal):

group like terms eg 4 x + 7x - 8 x = 3 x

OpenStudy (anonymous):

i remember learning this back in high school but @triciaal said it best.

OpenStudy (anonymous):

oh!! so easy.. i was totally over thinking it

OpenStudy (anonymous):

hs is so eh, lol

OpenStudy (anonymous):

yea i hated it to. the learning part anyways.

OpenStudy (anonymous):

ikr :(

OpenStudy (anonymous):

But when you graduate you can go to school for something useful in life. at least thats what i looked forward to.

OpenStudy (anonymous):

I'm going for engineering

OpenStudy (anonymous):

%80 of things i learned in high school is useless for what i wanted to do sadly high schools a "One size fits all" kind of learning.

OpenStudy (triciaal):

Math is always useful in life. Especially Algebra.

OpenStudy (anonymous):

Engineering a an great field. just expect a lot of math.

OpenStudy (anonymous):

is a*

OpenStudy (anonymous):

lol, i used to suck at algebra, to my surprise now, it's one of my best subjects. I'm just a bit sick and I have to catch up on missing work :( do you mind helping me more?

OpenStudy (triciaal):

yes engineering needs math and physics

OpenStudy (anonymous):

im sick aswell….from a flu shot .-., well if you have anything i understand yah sure here in car design you don't use a lot of math, at least for what i do. but yea give me a question

OpenStudy (anonymous):

1. Renting a bicycle at a local Bike shop costs $40 for the first day and $10 for each additional day. a. Create a chart of values and write an equation to find the total cost, C in terms of the number of days d.

OpenStudy (triciaal):

Hint y = m x + b m = 10 x= number of days and b = rental fee y = cost

OpenStudy (anonymous):

thanks :) you don't have to do the rest since it is my homework. Lol, and I don't want to cheat. So thanks xoxox

OpenStudy (anonymous):

that feel when you're trying to do 4 things at once xD

OpenStudy (anonymous):

ikr? lol